# Bangle functions are the generic basis for cluster algebras from punctured surfaces with boundary

### v2

@@ 5.1. Derksen–Weyman–Zelevinsky’s representation-theoretic Key Lemma
  • $h_{k}$ and $h_{k}^{\prime}$ are the $k^{th}$ components of the $\mathbf{h}$-vectors $\mathbf{h}_{\mathcal{M}}$ and $\mathbf{h}_{\overline{\mathcal{M}}}$, respectively, defined by [10, Equations (1.8) and (3.2)].
Moreover, the $\mathbf{g}$-vector $\mathbf{g}_{\mathcal{P}(Q,S)}^{\operatorname{DWZ}}(\mathcal{M})=(g_{1},\ldots,g_{n})$ defined by [10, Equation (1.13)] satisfies
@@
(5.3) $g_{j}^{\prime}:=\left\{\begin{array}[]{ll}-g_{k},&\text{if $j=k$};\\ \\ g_{j}+[b_{jk}]_{+}g_{k}-b_{jk}h_{k},&\text{if $j\neq k$}.\\ \end{array}\right.$
@@
By [10, Proposition 10.4 and Remark 10.8], $\mathbf{g}_{\mathcal{M}}^{\operatorname{DWZ}}=\mathbf{g}_{\mathcal{M}}^{\operatorname{inj}}$ (see also (3.2)).
@@ 5.2. Statement of the Combinatorial Key Lemma
(5.6) $g_{j}^{\prime}:=\left\{\begin{array}[]{ll}-g_{k},&\text{if $j=k$};\\ \\ g_{j}+[b_{jk}]_{+}g_{k}-b_{jk}h_{k},&\text{if $j\neq k$}.\\ \end{array}\right.$
@@ 5.5. Extendability of perfect matchings
Let $\mathcal{H}=[H_{1},\ldots,H_{s}]$ and $\mathcal{K}=[K_{1},\ldots,K_{t}]$ be snake graphs contained in a band graph $\mathcal{G}$, such that $H_{i}\neq K_{j}$ for all $i,j$. Let $P_{\mathcal{H}}$ and $P_{\mathcal{K}}$ be perfect matchings of $\mathcal{H}$ and $\mathcal{K}$, respectively. We say that $P_{\mathcal{H}}$ is extendable to $P_{\mathcal{K}}$ in $\mathcal{G}$ if there exists a perfect matching $P$ of the union $[H_{1},H_{2},\ldots,K_{t-1},K_{t}]$ such that the orientation induced by $P$ on the diagonals of $\mathcal{H}$ and $\mathcal{K}$ agrees with the orientation induced by $P_{\mathcal{H}}$ and $P_{\mathcal{K}}$, respectively.
@@
This follows from the existence of a minimal matching on $[H_{s},\ldots,K_{1}]$. ∎
@@
Case 4b): Here $b_{\gamma^{\prime}a}=2$ and $b_{\gamma^{\prime}b}=b_{\gamma^{\prime}c}=-1$. Taking $P_{i}$ such that diagonals $b$ and $c$ are negative, and all other diagonals are positive shows $h_{k}\leq 1-k$. Moreover, since there are only $k-1$ diagonals labelled by $\gamma^{\prime}$ then $h_{k}=1-k$.
@@
decomposes as the sum of $2^{m}$ collections of perfect matchings – each of these sums corresponding to a choice of orientation on each of the diagonals $d_{1},\ldots,d_{m}$. Some of these sums may be empty; by convention we define an empty sum to be 00. Specifically, if we let $(d_{i_{1}},\ldots,d_{i_{k}})$ denote the collection of good matchings which induce positive orientation on $d_{i_{1}},\ldots,d_{i_{k}}$ and negative orientation on the remaining diagonals, then we have the equality
@@ 5.7. Local $\mathbf{g}$-vectors
Moreover, the $k^{th}$ component of $g_{G(T,\alpha)}-\displaystyle\sum_{\beta\in\mathcal{L}_{k}}g_{G(T,\beta)}$ is 00.
@@
Furthermore, let $\gamma\in T$ be the arc corresponding to $k$. Note that $\gamma$ is an edge label of $G(T,\alpha)$ if and only if this edge belongs to $G(T,\beta)$ for some $\beta\in\mathcal{L}_{k}$. Consequently, the $k^{th}$ component of $g_{G(T,\alpha)}-\displaystyle\sum_{\beta\in\mathcal{L}_{k}}g_{G(T,\beta)}$ is 00. ∎
@@
Theorem 5.18 should be compared to [44, Corollary 6.15-(2)], though the reader is advised to be wary of signs and conventions. Our techniques provide a new, independent proof of Theorem 5.18.
@@ 6.1. The bangle function of a tagged arc belongs to the generic basis
Since the boundary of $\Sigma$ is not empty, by [15, Proposition 7.10] there exists a sequence of flips that transforms $T$ in a tagged triangulation $T_{\alpha}=f_{k_{m}}\cdots f_{k_{1}}(T)$ containing $\alpha$. The negative simple representation $\mathcal{S}_{\alpha}^{-}(A(T_{\alpha}))$ (see [9, the paragraph preceding Proposition 10.15] or [10, Equation (1.15)]) is the unique point in an irreducible component $Z_{T_{\alpha},\alpha}$ for the Jacobian algebra $A(T_{\alpha})$. This component $Z_{T_{\alpha},\alpha}$ is obviously $\tau^{-}$-rigid, i.e., it is a $\tau^{-}$-reduced component with $E$-invariant zero. By Proposition 3.8 and the invariance of the (injective) $E$-invariant under mutations [10, Theorem 7.1], $Z_{T,\alpha}:=\widetilde{\mu}_{k_{m}}\cdots\widetilde{\mu}_{k_{1}}(Z_{T_{\alpha},\alpha})$ is a $\tau^{-}$-reduced component with $E$-invariant zero for $A(T)$, and the generic value $CC_{A(T)}(Z_{T,\alpha})$ is given by $CC_{A(T)}(\mathcal{M}(T,\alpha))$, where $\mathcal{M}(T,\alpha):=\mu_{k_{m}}\cdots\mu_{k_{1}}(\mathcal{S}_{\alpha}^{-}(A(T_{\alpha})))$.
Now, by [19, Cor. 6.3] and [10, Eqn. (2.14) and Thm. 5.1], $CC_{A(T)}(\mathcal{M}(T,\alpha))$ is the cluster variable that corresponds to $\alpha$ according to [15, Theorem 7.11]. On the other hand, by [43, Theorems 4.9, 4.16 and 4.20], $CC_{A(T)}(\mathcal{M}(T,\alpha))=\operatorname{MSW}(G(T,\alpha))$. ∎
@@ 6.2. The bangle function of a closed curve belongs to the generic basis
is a generically $\tau^{-}$-reduced indecomposable irreducible component of the representation space $\operatorname{rep}(A(T_{\alpha}),\mathbf{d})$, where $\mathbf{d}:=\underline{\dim}(M(T_{\alpha},\alpha,\lambda,1))$, with
@@
Since the boundary of $\Sigma$ is non-empty, there exists a finite sequence $(T_{0},T_{1},\ldots,T_{m})$ of tagged triangulations, with $T_{0}=T_{\alpha}$ and $T_{m}=T$, such that for each $i=1,\ldots,n$, the triangulation $T_{i}$ is obtained from $T_{i-1}$ by flipping an arc $k_{i}\in T_{i-1}$. By [36, Theorem 8.1] and [9, Proposition 3.7], the Jacobian algebra $A(T):=\mathcal{P}_{\mathbb{C}}(Q^{\prime}(T),S^{\prime}(T))$ is isomorphic to the Jacobian algebra of $\mu_{k_{m}}\cdots\mu_{k_{1}}(Q^{\prime}(T_{\alpha}),S^{\prime}(T_{\alpha}))$. By Proposition 3.8,
@@
is a $\tau^{-}$-reduced indecomposable irreducible component of the representation varieties of $\mathcal{P}_{\mathbb{C}}(\mu_{k_{m}}\cdots\mu_{k_{1}}(Q^{\prime}(T_{\alpha}),S^{\prime}(T_{\alpha})))\cong A(T)$. By Theorem 3.12, we have $Z_{T,\alpha}=\pi_{T}(\alpha)$.
@@
At this point, we have shown that the identity $\mathbf{g}_{A(T^{\prime})}(Z_{T^{\prime},\alpha})=\mathbf{g}_{G(T,\alpha)}$ holds for every tagged triangulation, which allows us to see that Theorem 5.3 and [10, Theorem 5.1 and Lemma 5.2] and the particular equality $F_{M(T_{\alpha},\alpha,\lambda,1)}=F_{G(T_{\alpha},\alpha)}$ that was established in Proposition 4.1, imply that $F_{M(T,\alpha,\lambda,1)}=F_{G(T,\alpha)}$.
@@ 6.3. Main result
(a)(i) Note that the case when $L$ consists entirely of tagged arcs is known due to the work of Derksen-Weyman-Zelevinsky [10, Theorem 5.1]. The case when $L$ contains a closed curve follows immediately from Theorem 5.3, Proposition 4.1, and  [10, Lemma 5.2].
(a)(ii) By Theorem 3.12, we know that there is a bijection between the set of single laminates in $\operatorname{Lam}(\mathbf{\Sigma})$ and the set of indecomposable $\tau^{-}$-reduced components of the representation spaces of the Jacobian algebra $A(T)$. Furthermore, we have proved in Propositions 6.1 and 6.2 that for the indecomposable $\tau^{-}$-reduced component $Z_{T,\alpha}=\pi_{T}(\alpha)$ associated to each single laminate $\alpha$, the coefficient-free generic Caldero-Chapoton function $CC_{A(T)}(Z_{T,\alpha})$ is equal to $\operatorname{MSW}(G(T,\alpha))$.
On the other hand, it follows easily from [6, Lemma 4.11] and the definitions, that when $Z=\overline{Z_{1}\oplus Z_{2}}$ for $Z_{1},Z_{2}\in\operatorname{DecIrr}^{\tau^{-}}(A(T)$ (with $E_{A(T)}(Z_{1},Z_{2})=0$) then we have $CC_{A(T)}(Z)=CC_{A(T)}(Z_{1})\cdot CC_{A(T)}(Z_{2})$. Since $\pi_{T}$ is an isomorphism of (tame) partial KRS-monids by Theorem 3.12, our claim follows from these observations and the definition of $\operatorname{MSW}(T,L)$.
(b) Since $\pi_{T}$ is bijective we get by (a) in particular $\mathcal{B}^{\circ}(\mathbf{\Sigma},T)=\mathcal{B}_{A(T)}$. Next, suppose that $\mathbf{\Sigma}$ is not the dreaded torus $\mathbf{\Sigma}_{1}$ (i.e. we exclude the case $g(\Sigma)=1,\mathbb{P}=\emptyset,\mathbb{M}=\{m\}$). Then, since $\partial\Sigma\neq\emptyset$, by [21, Thm. 1.4] the potential $S^{\prime}(T)$ is up to right equivalence the unique non-degenerate potential for $Q^{\prime}(T)$. Thus, in this case $\mathcal{B}_{A(T)}$ is by [22, Cor. 6.14] a basis of the Caldero-Chapoton algebra $\mathcal{C}_{A(T)}$.
@@
(c) We have under this hypothesis $\mathcal{A}(\mathbf{\Sigma})=\mathcal{U}(\mathbf{\Sigma})$ either by the main result from [2], or because $\mathcal{A}(\mathbf{\Sigma})$ is locally acyclic by [41, Thm. 10.6 & Thm. 4.1]. So our claim follows from (b) together with the inclusions $\mathcal{A}(\mathbf{\Sigma})\subset\mathcal{C}_{A(T)}\subset\mathcal{U}(\mathbf{\Sigma})$. ∎
7. Concluding remarks and open problems
During the proofs of Theorems 6.1 and 6.2, we have shown that, given $\alpha$, one can find a (tagged) triangulation $T_{\alpha}$ for which one can further find a very concrete (decorated) representation $\mathcal{M}(T_{\alpha},\alpha)$ of the Jacobian algebra $A(T_{\alpha})=\mathcal{P}(Q(T_{\alpha}),S(T_{\alpha}))$ such that
@@ 7. Concluding remarks and open problems
Now, although $\mathcal{M}(T_{\alpha},\alpha)$ is a well-defined representation, we have not provided an explicit computation of $\mathcal{M}(T,\alpha)$ in general. Of course, when $\mathbb{P}=\varnothing$, the Jacobian algebra $A(T)$ is gentle, and $\mathcal{M}(T,\alpha)$ can be written down explicitly. When $\mathbb{P}\neq\varnothing$ but the signature of $T$ is zero, the Jacobian algebra $A(T)$ is skewed-gentle, and $\mathcal{M}(T,\alpha)$ can be computed explicitly as well, cf. [8, 20, 24]. But when $\mathbb{P}\neq\varnothing$ and $T$ is arbitrary, the decorated representation $\mathcal{M}(T,\alpha)$ still remains to be explicitly computed in general. It should be noticed that the naive candidate for $\mathcal{M}(T,\alpha)$, namely, the obvious string or band representation of the quiver $Q(T)$ induced by $\alpha$, typically fails to be annihilated by the cyclic derivatives of the potential $S(T)$, see e.g. [34, Example 6.2.7]. Explicit computations of $\mathcal{M}(T,\alpha)$ have been carried out by the second named author in [34] in the following situations:

### v3

@@ 5.1. Derksen–Weyman–Zelevinsky’s representation-theoretic Key Lemma
  • $h_{k}$ and $h_{k}^{\prime}$ are the $k^{th}$ components of the $\mathbf{h}$-vectors $\mathbf{h}_{\mathcal{M}}$ and $\mathbf{h}_{\overline{\mathcal{M}}}$, respectively, defined by [11, Equations (1.8) and (3.2)].
Moreover, the $\mathbf{g}$-vector $\mathbf{g}_{\mathcal{P}(Q,S)}^{\operatorname{DWZ}}(\mathcal{M})=(g_{1},\ldots,g_{n})$ defined by [11, Equation (1.13)] satisfies
@@
[…]
@@
By [11, Proposition 10.4 and Remark 10.8], $\mathbf{g}_{\mathcal{M}}^{\operatorname{DWZ}}=\mathbf{g}_{\mathcal{M}}^{\operatorname{inj}}$ (see also (3.2)).
@@ 5.2. Statement of the Combinatorial Key Lemma
[…]
@@ 5.5. Extendability of perfect matchings
Let $\mathcal{H}=[H_{1},\ldots,H_{s}]$ and $\mathcal{K}=[K_{1},\ldots,K_{t}]$ be snake graphs contained in a band graph $\mathcal{G}$, such that $H_{i}\neq K_{j}$ for all $i,j$. Let $P_{\mathcal{H}}$ and $P_{\mathcal{K}}$ be perfect matchings of $\mathcal{H}$ and $\mathcal{K}$, respectively. We say that $P_{\mathcal{H}}$ is extendable to $P_{\mathcal{K}}$ in $\mathcal{G}$ if there exists a perfect matching $P$ of the union $H_{1},H_{2},\ldots,K_{t-1},K_{t}]$ such that the orientation induced by $P$ on the diagonals of $\mathcal{H}$ and $\mathcal{K}$ agrees with the orientation induced by $P_{\mathcal{H}}$ and $P_{\mathcal{K}}$, respectively.
@@
This follows from the existence of a minimal matching on $H_{s},\ldots,K_{1}]$. ∎
@@
Case 4b): Here $b_{\gamma^{\prime}a}=2$ and $b_{\gamma^{\prime}b}=b_{\gamma^{\prime}c}=-1$. Taking $P_{i}$ such that diagonals $b$ and $c$ are negative, and all other diagonals are positive shows $h_{k}\leq 1-k$. Moreover, since there are only k1k-1 diagonals labelled by $\gamma^{\prime}$ then $h_{k}=1-k$.
@@
decomposes as the sum of $2^{m}$ collections of perfect matchings – each of these sums corresponding to a choice of orientation on each of the diagonals $d_{1},\ldots,d_{m}$. Some of these sums may be empty; by convention we define an empty sum to be 0. Specifically, if we let $(d_{i_{1}},\ldots,d_{i_{k}})$ denote the collection of good matchings which induce positive orientation on $d_{i_{1}},\ldots,d_{i_{k}}$ and negative orientation on the remaining diagonals, then we have the equality
@@ 5.7. Local $\mathbf{g}$-vectors
Moreover, the $k^{th}$ component of $g_{G(T,\alpha)}-\displaystyle\sum_{\beta\in\mathcal{L}_{k}}g_{G(T,\beta)}$ is 0.
@@
Furthermore, let $\gamma\in T$ be the arc corresponding to $k$. Note that $\gamma$ is an edge label of $G(T,\alpha)$ if and only if this edge belongs to $G(T,\beta)$ for some $\beta\in\mathcal{L}_{k}$. Consequently, the $k^{th}$ component of $g_{G(T,\alpha)}-\displaystyle\sum_{\beta\in\mathcal{L}_{k}}g_{G(T,\beta)}$ is 0. ∎
@@
Theorem 5.18 should be compared to [45, Corollary 6.15-(2)], though the reader is advised to be wary of signs and conventions. Our techniques provide a new, independent proof of Theorem 5.18.
@@ 6.1. The bangle function of a tagged arc belongs to the generic basis
Since the boundary of $\Sigma$ is not empty, by [16, Proposition 7.10] there exists a sequence of flips that transforms $T$ in a tagged triangulation $T_{\alpha}=f_{k_{m}}\cdots f_{k_{1}}(T)$ containing $\alpha$. The negative simple representation $\mathcal{S}_{\alpha}^{-}(A(T_{\alpha}))$ (see [10, the paragraph preceding Proposition 10.15] or [11, Equation (1.15)]) is the unique point in an irreducible component $Z_{T_{\alpha},\alpha}$ for the Jacobian algebra $A(T_{\alpha})$. This component $Z_{T_{\alpha},\alpha}$ is obviously $\tau^{-}$-rigid, i.e., it is a $\tau^{-}$-regular component with $E$-invariant zero. By Proposition 3.8 and the invariance of the (injective) $E$-invariant under mutations [11, Theorem 7.1], $Z_{T,\alpha}:=\widetilde{\mu}_{k_{m}}\cdots\widetilde{\mu}_{k_{1}}(Z_{T_{\alpha},\alpha})$ is a $\tau^{-}$-regular component with $E$-invariant zero for $A(T)$, and the generic value $CC_{A(T)}(Z_{T,\alpha})$ is given by $CC_{A(T)}(\mathcal{M}(T,\alpha))$, where $\mathcal{M}(T,\alpha):=\mu_{k_{m}}\cdots\mu_{k_{1}}(\mathcal{S}_{\alpha}^{-}(A(T_{\alpha})))$.
Now, by [20, Cor. 6.3] and [11, Eqn. (2.14) and Thm. 5.1], $CC_{A(T)}(\mathcal{M}(T,\alpha))$ is the cluster variable that corresponds to $\alpha$ according to [16, Theorem 7.11]. On the other hand, by [44, Theorems 4.9, 4.16 and 4.20], $CC_{A(T)}(\mathcal{M}(T,\alpha))=\operatorname{MSW}(G(T,\alpha))$. ∎
@@ 6.2. The bangle function of a closed curve belongs to the generic basis
is a generically $\tau^{-}$-regular indecomposable irreducible component of the representation space $\operatorname{rep}(A(T_{\alpha}),\mathbf{d})$, where $\mathbf{d}:=\underline{\dim}(M(T_{\alpha},\alpha,\lambda,1))$, with
@@
Since the boundary of $\Sigma$ is non-empty, there exists a finite sequence $(T_{0},T_{1},\ldots,T_{m})$ of tagged triangulations, with $T_{0}=T_{\alpha}$ and $T_{m}=T$, such that for each $i=1,\ldots,n$, the triangulation $T_{i}$ is obtained from $T_{i-1}$ by flipping an arc $k_{i}\in T_{i-1}$. By [37, Theorem 8.1] and [10, Proposition 3.7], the Jacobian algebra $A(T):=\mathcal{P}_{\mathbb{C}}(Q^{\prime}(T),S^{\prime}(T))$ is isomorphic to the Jacobian algebra of $\mu_{k_{m}}\cdots\mu_{k_{1}}(Q^{\prime}(T_{\alpha}),S^{\prime}(T_{\alpha}))$. By Proposition 3.8,
@@
is a $\tau^{-}$-regular indecomposable irreducible component of the representation varieties of $\mathcal{P}_{\mathbb{C}}(\mu_{k_{m}}\cdots\mu_{k_{1}}(Q^{\prime}(T_{\alpha}),S^{\prime}(T_{\alpha})))\cong A(T)$. By Theorem 3.12, we have $Z_{T,\alpha}=\pi_{T}(\alpha)$.
@@
At this point, we have shown that the identity $\mathbf{g}_{A(T^{\prime})}(Z_{T^{\prime},\alpha})=\mathbf{g}_{G(T,\alpha)}$ holds for every tagged triangulation, which allows us to see that Theorem 5.3 and [11, Theorem 5.1 and Lemma 5.2] and the particular equality $F_{M(T_{\alpha},\alpha,\lambda,1)}=F_{G(T_{\alpha},\alpha)}$ that was established in Proposition 4.1, imply that $F_{M(T,\alpha,\lambda,1)}=F_{G(T,\alpha)}$.
@@ 6.3. Main result
(a)(i) Note that the case when $L$ consists entirely of tagged arcs is known due to the work of Derksen-Weyman-Zelevinsky [11, Theorem 5.1]. The case when $L$ contains a closed curve follows immediately from Theorem 5.3, Proposition 4.1, and  [11, Lemma 5.2].
(a)(ii) By Theorem 3.12, we know that there is a bijection between the set of single laminates in $\operatorname{Lam}(\mathbf{\Sigma})$ and the set of indecomposable $\tau^{-}$-regular components of the representation spaces of the Jacobian algebra $A(T)$. Furthermore, we have proved in Propositions 6.1 and 6.2 that for the indecomposable $\tau^{-}$-regular component $Z_{T,\alpha}=\pi_{T}(\alpha)$ associated to each single laminate $\alpha$, the coefficient-free generic Caldero-Chapoton function $CC_{A(T)}(Z_{T,\alpha})$ is equal to $\operatorname{MSW}(G(T,\alpha))$.
On the other hand, it follows easily from [7, Lemma 4.11] and the definitions, that when $Z=\overline{Z_{1}\oplus Z_{2}}$ for $Z_{1},Z_{2}\in\operatorname{DecIrr}^{\tau^{-}}(A(T)$ (with $E_{A(T)}(Z_{1},Z_{2})=0$) then we have $CC_{A(T)}(Z)=CC_{A(T)}(Z_{1})\cdot CC_{A(T)}(Z_{2})$. Since $\pi_{T}$ is an isomorphism of (tame) partial KRS-monids by Theorem 3.12, our claim follows from these observations and the definition of $\operatorname{MSW}(T,L)$.
(b) Since $\pi_{T}$ is bijective we get by (a) in particular $\mathcal{B}^{\circ}(\mathbf{\Sigma},T)=\mathcal{B}_{A(T)}$. Next, suppose that $\mathbf{\Sigma}$ is not the dreaded torus $\mathbf{\Sigma}_{1}$ (i.e. we exclude the case $g(\Sigma)=1,\mathbb{P}=\emptyset,\mathbb{M}=\{m\}$). Then, since $\partial\Sigma\neq\emptyset$, by [22, Thm. 1.4] the potential $S^{\prime}(T)$ is up to right equivalence the unique non-degenerate potential for $Q^{\prime}(T)$. Thus, in this case $\mathcal{B}_{A(T)}$ is by [23, Cor. 6.14] a basis of the Caldero-Chapoton algebra $\mathcal{C}_{A(T)}$.
@@
(c) We have under this hypothesis $\mathcal{A}(\mathbf{\Sigma})=\mathcal{U}(\mathbf{\Sigma})$ either by the main result from [3], or because $\mathcal{A}(\mathbf{\Sigma})$ is locally acyclic by [42, Thm. 10.6 & Thm. 4.1]. So our claim follows from (b) together with the inclusions $\mathcal{A}(\mathbf{\Sigma})\subset\mathcal{C}_{A(T)}\subset\mathcal{U}(\mathbf{\Sigma})$. ∎
7. Applications and open problems
During the proofs of Theorems 6.1 and 6.2, we have shown that, given any lamination $\alpha\in\operatorname{Lam}(\mathbf{\Sigma})$, one can find a (tagged) triangulation $T_{\alpha}$ and a very concretely constructed (decorated) representation $\mathcal{M}(T_{\alpha},\alpha)$ of the Jacobian algebra $A(T_{\alpha})=\mathcal{P}(Q(T_{\alpha}),S(T_{\alpha}))$ such that
@@ 7. Applications and open problems
As a consequence, we get the following immediate application.
Corollary 7.1.
Let $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ be a surface with marked points such that $\partial\Sigma\neq\emptyset$, and let $T$ be a tagged triangulation of $\mathbf{\Sigma}$.
Then for all laminations $\alpha\in\operatorname{Lam}(\mathbf{\Sigma})$, if $L$ is a closed curve (resp. a tagged arc) then the Euler-Poincaré characteristic $\chi(Gr_{e}(\mathcal{M}(T,\alpha)))$ equals the number of good matchings $P$ of the band graph (resp. loop graph) $G(T,\alpha)$ such that $y(P)=\prod_{i=1}^{n}y_{i}^{e_{i}}.$
Our ability to obtain the above result is quite peculiar. Note that, although $\mathcal{M}(T_{\alpha},\alpha)$ is a well-defined representation, we have not provided an explicit computation of $\mathcal{M}(T,\alpha)$ in general, but are still able to compute its Euler-Poincaré characteristic. Of course, when $\mathbb{P}=\varnothing$, the Jacobian algebra $A(T)$ is gentle, and $\mathcal{M}(T,\alpha)$ can be written down explicitly. When $\mathbb{P}\neq\varnothing$ but the signature of $T$ is zero, the Jacobian algebra $A(T)$ is skewed-gentle, and $\mathcal{M}(T,\alpha)$ can be computed explicitly as well, cf. [9, 21, 25]. However, when $\mathbb{P}\neq\varnothing$ and $T$ is arbitrary, the decorated representation $\mathcal{M}(T,\alpha)$ still remains to be explicitly computed in general. It should be noticed that the naive candidate for $\mathcal{M}(T,\alpha)$, namely, the obvious string or band representation of the quiver $Q(T)$ induced by $\alpha$, typically fails to be annihilated by the cyclic derivatives of the potential $S(T)$, see e.g. [35, Example 6.2.7]. Explicit computations of $\mathcal{M}(T,\alpha)$ have been carried out by the second named author in [35] in the following situations: